# for the Hall C SIDIS Collaboration

### v1

@@ Abstract
Semi-inclusive deep-inelastic lepton-nucleon scattering ($lN\rightarrow l^{{}^{\prime}}hX$) is an excellent tool to study the quark hadronization mechanism described by fragmentation functions (FF) Field and Feynman (1977). These FF describe how the quarks and gluons transform into color-neutral hadrons or photons during high-energy (hard) scattering processes. Pion semi-inclusive deep-inelastic scattering (SIDIS) is one such scattering process that allows access to the FF associated with the pions identified in the final state. As FF are intrinsically linked to confinement in quantum chromodynamics (QCD), studies of FF are critical for a complete understanding of the basic properties of QCD. Further, FF are the non-perturbative ingredient of the QCD factorization theorems Collins et al. (1988) used to analyze hard scattering processes and thereby provide insight into fundamental soft QCD quantities Albino (2010). The current knowledge of pion FF is based on global QCD analyses hkns; deFlorian et al. (2007, 2015); Albino et al. (2008); Moffat et al. (2021); Bacchetta et al. (2022) that are dominated by measurements from inclusive electron-positron ($e^{+}e^{-}$) annihilation into charged pions at very high energy scales (center-of-mass energy $>$ 10 GeV). Inclusive $e^{+}e^{-}$ annihilation is a clean process to study FF since it is independent of the parton distribution functions (PDF). However, it cannot distinguish between the light quark flavors or the quark and anti-quark FF. Thus, it cannot provide information about possible flavor dependence of FF – essential for a complete picture of FF as well as the spin structure of nucleons, in particular the transverse spin structure Signori et al. (2013). One of the most important advantages of SIDIS is the ability to constrain the flavor of the quark involved in the scattering process. Consequently, measuring the SIDIS process on protons and deuterons allows an independent extraction of the flavor dependence of FF. The SIDIS experiments conducted over the last decade have convincingly established that the collinear picture of the quark-parton model is too simple, highlighting the importance of the transverse structure of the hadrons. The flavor structure of FF is important to understand the flavor dependence of the transverse-momentum-dependent (TMD) FF Signori et al. (2013), and the relative differences between the observed single spin asymmetries of pions and kaons Airapetian et al. (2009); Adolph et al. (2015). Thus, SIDIS measurements provide a unique capability to study the flavor structure of FF at an energy scale that is complementary to that of $e^{+}e^{-}$ annihilation.
@@
We define the measured multiplicities for $\pi^{+}$ and $\pi^{-}$ production from hydrogen (p) and deuterium (d) targets, $M^{\pi^{\pm}}_{\mbox{p/d}}(x,Q^{2},z)$, as the ratio of the respective SIDIS cross section to the DIS cross section (see Eq. S1 in the online Supplementary Material). At leading order, assuming CS, and no difference in the $p_{T}$ dependence of the measured multiplicities, we can write two simple ratios,
@@
Most global analyses to extract PDF assume IS and CS in the PDF, which reduces the number of independent PDF by half. If we assume CS in the PDF but allow for non-zero CSV in FF, the multiplicity $M^{\pi^{\pm}}_{\mbox{p/d}}(x,Q^{2},z)$ for each target (H/D) and charged pion type can be written in terms of two favored FF, $D_{u\pi^{+}}(z)$, $D_{d\pi^{-}}(z)$, and two un-favored FF, $D_{d\pi^{+}}(z)$, $D_{u\pi^{-}}(z)$, respectively (see Eq. LABEL:eq:s1 in the Supplementary Material). Any difference between the two favored and the two un-favored FF is an indication of CSV in FF. The degree of CSV in the favored and un-favored FF can be quantified in terms of two parameters defined as:
@@
The experiment was carried out in the Fall of 2018 and the Spring of 2019, in Hall C at JLab. The experiment used the quasi-continuous wave electron beam with beam energies of 10.2 and 10.6 GeV and beam currents ranging from 2 $\mu$A to 70 $\mu$A. Additional details of the experiment are described in Sec. S-I of the Supplementary Material. The experimental yields were obtained from selected electron-pion coincidence events per milli-Coulomb of electrons incident on 1H, and 2H targets. The selected events passed cuts on momentum, scattering angles, and missing mass of the residual system, $M_{X}$, where $M_{X}$ was restricted to be above the resonance region ($M_{X}>$ 1.6 $\mathrm{GeV}/\mathrm{c}^{2}$). The yields were integrated over azimuthal angle ($\phi$) and $p_{T}$. The backgrounds from the target’s aluminum windows and accidental coincidences were subtracted. This normalized SIDIS pion electroproduction yield was corrected for all known inefficiencies of the two spectrometers such as the detector efficiencies (97%–99%), trigger efficiency (98%-99%), tracking efficiencies, computer and electronic live times (94%–99%). The corrected yields were binned in $z$ for the 8 different kinematic settings where the $x$ ranged from 0.3 to 0.6, $Q^{2}$ ranged from 3.1 to 5.5 GeV2 and the center-of-mass energy, $W$, ranged from 2.2 to 3.2 GeV. The table of kinematic settings is shown in Table  S1 of the Supplementary Material.
@@
A Monte Carlo (MC) simulation https://hallcweb.jlab.org/wiki/index.php/SIMC˙Monte˙Carlo of the SIDIS process was performed with the factorized form shown in Eq. 1. The CTEQ5 next-to-leading-order (NLO) PDF were used to parametrize $q(x,Q^{2})$ Lai et al. (1997) along with a parametrization of FF from global fits of SIDIS data Bosted and Christy (2008). The MC was used to smear parameterized PDF and FF over the experimental acceptance. The MC included corrections due to radiative tails from exclusive pion electroproduction, pion decay, and electroproduction of $\rho^{0}$ mesons and $\Delta$(1232) resonances. Additional details about the models used in the simulation can be found in Refs. Bhatt ; Jia . The Monte Carlo yields were integrated over the same phase space as the measured yields.
@@
The corrected experimental yields for the 8 different settings were used along with the Monte Carlo yield and the model cross section to obtain the four multiplicities, $M^{\pi^{\pm}}_{\mbox{p/d}}(x,Q^{2},z)$. The systematic uncertainty of the extracted multiplicities is 2.8%, it is listed in Table S2 and described in the Supplementary Material. The four multiplicities as a function of $W^{2}$ are shown in Fig. S7 of the Supplementary Material for 8 different $z$ bins ranging from $z=$ 0.3 to 0.7. Note that our data confirm that the $p_{T}$ dependence for the $\pi^{\pm}$ multiplicities from hydrogen and deuterium are identical within the small $p_{T}$ range covered. The four multiplicities were used to form the two ratios $R_{1}(z)$ and $R_{2}(z)$, and were also used to obtain four FF ($D_{u\pi^{+}}(z),D_{d\pi^{-}}(z)$ and $D_{u\pi^{-}}(z),D_{d\pi^{+}}(z)$) by simultaneously solving a system of four equations (Eq. LABEL:eq:s1 in the Supplementary Material).
@@
The extracted sum and difference ratios and their statistical uncertainties are shown as a function of $z$ in Fig. 1. For the sum and difference ratios, many systematic uncertainties cancel to first order resulting in a net 2.2% systematic uncertainty shown by the magenta band in the bottom-right panel (see the Supplemental Material for additional details). The open (closed) symbols are with (without) subtraction of the diffractive $\rho^{0}$ contributions. The negligible difference between them shows that the diffractive $\rho^{0}$ contribution to the pion yield has very little impact on these ratios. The solid (dotted) lines are the expectations for models with CS/IS. The dot-dashed and dashed curves use FF from the global fits by the DSS deFlorian et al. (2007, 2015) and MAP Cerutti et al. (2023) collaborations, respectively. The uncertainty for the DSS curves is not shown because, unlike the experiment and MAP results, they are integrated over all $p_{T}$. At the highest $W$ (3.2 GeV) the two ratios are remarkably independent of $z$ over the entire range ($z$ = 0.3 - 0.7) and are also consistent with the magnitude predicted by the global fits to existing data. In other words, the results agree with the CS/IS expectation. The sum ratio $R_{1}$ slowly but steadily deviates from the CS expectation with decreasing $W$ (increasing $x$), both in terms of the $z$ independence and the magnitude. Similarly, the difference ratio also shows increasingly large deviations from the CS expectation with decreasing $W$. These deviations may indicate the importance of higher twist contributions to the SIDIS cross sections at low $W$. These results also indicate that even for the limited range of $p_{T}$ covered in this experiment, CS/IS seems to be valid for $W>$ 3 GeV. Moreover, the sum/difference ratio from the previous JLab 6 GeV experiment Navasardyan et al. (2007) (shown as black triangles in the bottom left panel) agrees remarkably well with the current results. These older ratios were obtained at the same $x$ = 0.32, but at significantly lower $W$ and $Q^{2}$ of 2.4 GeV and 2.3 GeV2 respectively. This seems to indicate that $x$ may also be relevant for tests of CS/IS.
@@
The four multiplicities were also used to extract four FF which are shown as a function of $z$ in Fig. 2 for the highest $W$ (lowest $x$) setting ($W$= 3.2 GeV, $x$=0.3). Note that the JAM collaboration (solid lines) assumes CS/IS for all FF, while the DSS collaboration (dashed lines) assumes CS/IS only for the un-favored FF. The variation of the extracted FF due to the normalization type uncertainties of the multiplicities was used to determine the systematic uncertainty of the FF. The extracted FF for all 8 settings are shown in Fig. S8 of the Supplementary Material. The two favored and two un-favored FF were used to form the favored and un-favored $\delta_{CSV}(z)$ parameters as defined in Eq. 4 and are shown in Fig. 3. The settings with similar $W$ and $x$ were averaged to reduce the 8 kinematic settings to 4. The variation in the $\delta_{CSV}$ parameter due to the choice of PDF and normalization type uncertainty was used to determine the systematic uncertainties of $\delta_{CSV}$. The systematic uncertainty is shown by the shaded bands. The favored $\delta_{CSV}$ parameter is essentially zero within the experimental uncertainties over the entire range of $z$ and $W$. They are also consistent with the expectations of the global fits by DSS deFlorian et al. (2015) and Peng and Ma Peng and Ma (2023) but not with the fits by the MAP collaboration Cerutti et al. (2023).
@@
$x$ $Q^{2}$ $W$ $z$ $M^{\pi^{+}}_{p}$ $\delta M^{\pi^{+}}_{p}$ $M^{\pi^{+}}_{d}$ $\delta M^{\pi^{+}}_{d}$ $M^{\pi^{-}}_{p}$ $\delta M^{\pi^{-}}_{p}$ $M^{\pi^{-}}_{d}$ $\delta M^{\pi^{-}}_{\ d}$
(GeV2) (GeV)

### v2

@@ Abstract
Semi-inclusive deep-inelastic lepton-nucleon scattering ($lN\rightarrow l^{{}^{\prime}}hX$) is an excellent tool to study the quark hadronization mechanism described by fragmentation functions (FF) Field and Feynman (1977). These FF describe how the quarks and gluons transform into color-neutral hadrons or photons during high-energy (hard) scattering processes. Pion semi-inclusive deep-inelastic scattering (SIDIS) is one such scattering process that allows access to the FF associated with the pions identified in the final state. As FF are intrinsically linked to confinement in quantum chromodynamics (QCD), studies of FF are critical for a complete understanding of the basic properties of QCD. Further, FF are the non-perturbative ingredient of the QCD factorization theorems Collins et al. (1988) used to analyze hard scattering processes and thereby provide insight into fundamental soft QCD quantities Albino (2010). The current knowledge of pion FF is based on global QCD analyses Hirai et al. (2007); deFlorian et al. (2007, 2015); Albino et al. (2008); Moffat et al. (2021); Bacchetta et al. (2022) that are dominated by measurements from inclusive electron-positron ($e^{+}e^{-}$) annihilation into charged pions at very high energy scales (center-of-mass energy $>$ 10 GeV). Inclusive $e^{+}e^{-}$ annihilation is a clean process to study FF since it is independent of the parton distribution functions (PDF). However, it cannot distinguish between the light quark flavors or the quark and anti-quark FF. Thus, it cannot provide information about possible flavor dependence of FF – essential for a complete picture of FF as well as the spin structure of nucleons, in particular the transverse spin structure Signori et al. (2013). One of the most important advantages of SIDIS is the ability to constrain the flavor of the quark involved in the scattering process. Consequently, measuring the SIDIS process on protons and deuterons allows an independent extraction of the flavor dependence of FF. The SIDIS experiments conducted over the last decade have convincingly established that the collinear picture of the quark-parton model is too simple, highlighting the importance of the transverse structure of the hadrons. The flavor structure of FF is important to understand the flavor dependence of the transverse-momentum-dependent (TMD) FF Signori et al. (2013), and the relative differences between the observed single spin asymmetries of pions and kaons Airapetian et al. (2009); Adolph et al. (2015). Thus, SIDIS measurements provide a unique capability to study the flavor structure of FF at an energy scale that is complementary to that of $e^{+}e^{-}$ annihilation.
@@
We define the measured multiplicities for $\pi^{+}$ and $\pi^{-}$ production from hydrogen (p) and deuterium (d) targets, $M^{\pi^{\pm}}_{\mbox{p/d}}(x,Q^{2},z)$, as the ratio of the respective SIDIS cross section to the DIS cross section (see Eq. S1 in the online Supplementary Material sup ). At leading order, assuming CS, and no difference in the $p_{T}$ dependence of the measured multiplicities, we can write two simple ratios,
@@
Most global analyses to extract PDF assume IS and CS in the PDF, which reduces the number of independent PDF by half. If we assume CS in the PDF but allow for non-zero CSV in FF, the multiplicity $M^{\pi^{\pm}}_{\mbox{p/d}}(x,Q^{2},z)$ for each target (H/D) and charged pion type can be written in terms of two favored FF, $D_{u\pi^{+}}(z)$, $D_{d\pi^{-}}(z)$, and two un-favored FF, $D_{d\pi^{+}}(z)$, $D_{u\pi^{-}}(z)$, respectively (see Eq. LABEL:eq:s1 in the Supplementary Material sup ). Any difference between the two favored and the two un-favored FF is an indication of CSV in FF. The degree of CSV in the favored and un-favored FF can be quantified in terms of two parameters defined as:
@@
The experiment was carried out in the Fall of 2018 and the Spring of 2019, in Hall C at JLab. The experiment used the quasi-continuous wave electron beam with beam energies of 10.2 and 10.6 GeV and beam currents ranging from 2 $\mu$A to 70 $\mu$A. Additional details of the experiment are described in Sec. S-I of the Supplementary Material sup . The experimental yields were obtained from selected electron-pion coincidence events per milli-Coulomb of electrons incident on 1H, and 2H targets. The selected events passed cuts on momentum, scattering angles, and missing mass of the residual system, $M_{X}$, where $M_{X}$ was restricted to be above the resonance region ($M_{X}>$ 1.6 $\mathrm{GeV}/\mathrm{c}^{2}$). The yields were integrated over azimuthal angle ($\phi$) and $p_{T}$. The backgrounds from the target’s aluminum windows and accidental coincidences were subtracted. This normalized SIDIS pion electroproduction yield was corrected for all known inefficiencies of the two spectrometers such as the detector efficiencies (97%–99%), trigger efficiency (98%-99%), tracking efficiencies, computer and electronic live times (94%–99%). The corrected yields were binned in $z$ for the 8 different kinematic settings where the $x$ ranged from 0.3 to 0.6, $Q^{2}$ ranged from 3.1 to 5.5 GeV2 and the center-of-mass energy, $W$, ranged from 2.2 to 3.2 GeV. The table of kinematic settings is shown in Table  S1 of the Supplementary Material sup .
@@
A Monte Carlo (MC) simulation https://hallcweb.jlab.org/wiki/index.php/SIMC˙Monte˙Carlo of the SIDIS process was performed with the factorized form shown in Eq. 1. The CTEQ5 next-to-leading-order (NLO) PDF were used to parametrize $q(x,Q^{2})$ Lai et al. (1997) along with a parametrization of FF from global fits of SIDIS data Bosted and Christy (2008). The MC was used to smear parameterized PDF and FF over the experimental acceptance. The MC included corrections due to radiative tails from exclusive pion electroproduction, pion decay, and electroproduction of $\rho^{0}$ mesons and $\Delta$(1232) resonances. Additional details about the models used in the simulation can be found in Refs. Bhatt ; Jia ; Bosted et al. . The Monte Carlo yields were integrated over the same phase space as the measured yields.
@@
The corrected experimental yields for the 8 different settings were used along with the Monte Carlo yield and the model cross section to obtain the four multiplicities, $M^{\pi^{\pm}}_{\mbox{p/d}}(x,Q^{2},z)$. The systematic uncertainty of the extracted multiplicities is 2.8%, it is listed in Table S2 and described in the Supplementary Material sup . The four multiplicities as a function of $W^{2}$ are shown in Fig. S7 of the Supplementary Material sup for 8 different $z$ bins ranging from $z=$ 0.3 to 0.7. Note that our data confirm that the $p_{T}$ dependence for the $\pi^{\pm}$ multiplicities from hydrogen and deuterium are identical within the small $p_{T}$ range covered. The four multiplicities were used to form the two ratios $R_{1}(z)$ and $R_{2}(z)$, and were also used to obtain four FF ($D_{u\pi^{+}}(z),D_{d\pi^{-}}(z)$ and $D_{u\pi^{-}}(z),D_{d\pi^{+}}(z)$) by simultaneously solving a system of four equations (Eq. LABEL:eq:s1 in the Supplementary Material sup ).
@@
The extracted sum and difference ratios and their statistical uncertainties are shown as a function of $z$ in Fig. 1. For the sum and difference ratios, many systematic uncertainties cancel to first order resulting in a net 2.2% systematic uncertainty shown by the magenta band in the bottom-right panel (see the Supplemental Material sup for additional details). The open (closed) symbols are with (without) subtraction of the diffractive $\rho^{0}$ contributions. The negligible difference between them shows that the diffractive $\rho^{0}$ contribution to the pion yield has very little impact on these ratios. The solid (dotted) lines are the expectations for models with CS/IS. The dot-dashed and dashed curves use FF from the global fits by the DSS deFlorian et al. (2007, 2015) and MAP Cerutti et al. (2023) collaborations, respectively. The uncertainty for the DSS curves is not shown because, unlike the experiment and MAP results, they are integrated over all $p_{T}$. At the highest $W$ (3.2 GeV) the two ratios are remarkably independent of $z$ over the entire range ($z$ = 0.3 - 0.7) and are also consistent with the magnitude predicted by the global fits to existing data. In other words, the results agree with the CS/IS expectation. The sum ratio $R_{1}$ slowly but steadily deviates from the CS expectation with decreasing $W$ (increasing $x$), both in terms of the $z$ independence and the magnitude. Similarly, the difference ratio also shows increasingly large deviations from the CS expectation with decreasing $W$. These deviations may indicate the importance of higher twist contributions to the SIDIS cross sections at low $W$. These results also indicate that even for the limited range of $p_{T}$ covered in this experiment, CS/IS seems to be valid for $W>$ 3 GeV. Moreover, the sum/difference ratio from the previous JLab 6 GeV experiment Navasardyan et al. (2007) (shown as black triangles in the bottom left panel) agrees remarkably well with the current results. These older ratios were obtained at the same $x$ = 0.32, but at significantly lower $W$ and $Q^{2}$ of 2.4 GeV and 2.3 GeV2 respectively. This seems to indicate that $x$ may also be relevant for tests of CS/IS.
@@
The four multiplicities were also used to extract four FF which are shown as a function of $z$ in Fig. 2 for the highest $W$ (lowest $x$) setting ($W$= 3.2 GeV, $x$=0.3). Note that the JAM collaboration (solid lines) assumes CS/IS for all FF, while the DSS collaboration (dashed lines) assumes CS/IS only for the un-favored FF. The variation of the extracted FF due to the normalization type uncertainties of the multiplicities was used to determine the systematic uncertainty of the FF. The extracted FF for all 8 settings are shown in Fig. S8 of the Supplementary Material sup . The two favored and two un-favored FF were used to form the favored and un-favored $\delta_{CSV}(z)$ parameters as defined in Eq. 4 and are shown in Fig. 3. The settings with similar $W$ and $x$ were averaged to reduce the 8 kinematic settings to 4. The variation in the $\delta_{CSV}$ parameter due to the choice of PDF and normalization type uncertainty was used to determine the systematic uncertainties of $\delta_{CSV}$. The systematic uncertainty is shown by the shaded bands. The favored $\delta_{CSV}$ parameter is essentially zero within the experimental uncertainties over the entire range of $z$ and $W$. They are also consistent with the expectations of the global fits by DSS deFlorian et al. (2015) and Peng and Ma Peng and Ma (2023) but not with the fits by the MAP collaboration Cerutti et al. (2023).
@@
$x$ $Q^{2}$(GeV2) $W$ (GeV) $z$ $M^{\pi^{+}}_{p}$ $\delta M^{\pi^{+}}_{p}$ $M^{\pi^{+}}_{d}$ $\delta M^{\pi^{+}}_{d}$ $M^{\pi^{-}}_{p}$ $\delta M^{\pi^{-}}_{p}$ $M^{\pi^{-}}_{d}$ $\delta M^{\pi^{-}}_{\ d}$